Ted walked for 90 minutes and biked for 30 minutes. He bikes 4 times faster than he walks. If the total distance traveled is 10.5 miles. What was his walking speed in miles per hour?
Answer by checkley75(3666)

1.5x+.5*4x=10.5
1.5x+2x=10.5
3.5x=10.5
x=10.5/3.5
x=3 mph for the walking rate.
proof
1.5*3+.5*4*3=10.5
4.5+6=10.5
10.5=10.5